OpenStudy (anonymous):

Find x belongs to [0,pi/2] such that cos^2 x -3 sin^2 x = 0.

jhonyy9 (jhonyy9):

like a first step you know that cos^2 x +sin^2 x =1 so than from this sin^2 x = 1 -cos^2 x subtitute this in equation in place of sin^2 x and will get one equation with cos x like unknowed term so note cos x = t and will get an quadratic equation in t so what you need solve it and after you will get the roots of quadratic make them equal cos x and so will get the right result(s) hope this will help you good luck

OpenStudy (bradely):

cos^2 x -3 sin^2 x = 0. cos^2 x =3 sin^2 x sin^2 x /cos^2 x=1/3 tan^2x =1/3 tanx=1/sqrt3 tanx= tan30 x=npi+pi/3 x=30 Source: http://www.mathskey.com/question2answer/ask